Meeting Notes for 20 January 2010

Present: TomL, Phong, Bob, Mihael, Liz, TomJ
Absent: Marge


  • Bob's finished new documentation - should be pushed this Friday to production
  • Data blessing meeting 18 Feb.


  • LIGO data feed has resumed; we are working to get the data working in existing-Bluestone
  • Working on the Bluestone re-architecting


  • See meeting notes from last week
  • Milestones are needed for our 1 June rollout.

-- Main.PhongNguyen - 20 Jan 2010
Topic revision: r9 - 2010-01-27, PhongNguyen
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback